REMPI SPECTROSCOPY OF THE 3pπ^{2}Π RYDBERG STATE OF HCO

\underline{Hartmut G. Hedderich}, Eric E. Mayer, and Edward R. Grant, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393.

The (010), (020), (030), and (001) bands of the 3pπ^{2}Π Rydberg state of the formyl radical have been observed by resonant (1+1) photon ionization spectroscopy from the ^{2}A′ vibronic ground state. Subbands have been re-assigned considering the D_{∞h} symmetry of HCO in the 3pπ-state. Renner-Teller effect and Fermi resonances have been analyzed for different subbands.

Time required: 10 min Session in which paper is recomended for presentation: 2 Comment: